an isosceles right triangle has area of 98 cm to the second power ..find the length of each side
Answer by nycsharkman(136)

In a right triangle, one of the angles is a right angle—an angle of 90 degrees. A right triangle may be isosceles or scalene.
The formula to use is:
A = 1/2 times base times height
Let x = length of each side (base and height)
98 = 1/2(x)(x)
98 = 1/2(x^2)
Divide both sides by 1/2.
98 divided by 1/2 = x^2
196 = x^2
Take square root of both sides.
+ - 14 = x
Since we are looking for length of sides, we MUST REJECT a negative answer for x because distance CANNOT be negative.
So, x = 14.
This means that each side measures 14cm.
Did you follow?